# IMU: Influence-guided Machine Unlearning

### v2

@@ 5 Conclusion, Limitations, Broader Impacts
In this paper, we introduce the retain-data-free MU method IMU, a novel paradigm that dynamically allocates attention to each data point according to its estimated influence value. Extensive experiments on various unlearning settings across vision and language tasks demonstrate that IMU effectively strikes the balance between forget quality and model utility, showing superiority compared to existing baselines.
However, our method still requires estimating the influence function for each forgetting data point. As a result, the computational cost scales with the number of data points, which can be time-consuming in real-world scenarios. One possible alternative is to estimate the influence at the mini-batch level rather than per data point. While this may reduce computation time, it could also introduce estimation errors. Besides, adversarial examples or outliers may interfere with influence estimation, leading to inaccurate decisions.
As MU becomes increasingly important, our method offers a practical solution that does not rely on the retain set. However, care must be taken to ensure robustness against adversarial manipulation of influence scores and to mitigate any potential misuse of the method.
@@
6 Details
6.1 Implementation details
All the experiments were performed on NVIDIA RTX3090 with Intel Xeon processors. For image classification scenario, the original model and retrained model were both trained with over $182$ epochs using the SGD optimizer with a cosine-scheduled, and learning rate was initialized at $0.1$. For both NPO and RL, training spans $10$ epochs within the interval $10^{-4},10^{-1}]$, and $\beta$ was selected from the range $[0.1,5]$. GA’s training settings around a $5$-epoch learning rate search within the interval $10^{-5},10^{-2}]$. In the case of IU, we explored the parameter $\alpha$ within the range $[1,20]$. For SSD, searching for the parameter selection weighting $\alpha$ and dampening constant $\gamma$ was executed within the range $[1,100]$ and $[0.1,1]$, respectively, while searching for the learning rate within the range $10^{-4},10^{-2}]$. For the SCAR method, the value of temperature was defined within the range $[2,5]$, the parameter $\lambda_{2}$ was set as $0.01$, and $\lambda_{1}$ was in the range of $[1,10]$ with the learning rate searching in the interval $10^{-6},10^{-2}]$. Learning rate was selected in the range $10^{-4},10^{-2}]$ for SalUn and let sparsity ratio equal to $0.5$. Lastly, for IMU method, we trained for 3 epochs, searching for learning rates in the range $10^{-5},10^{-2}]$ and $l_{1}$ regularization intensity in the range $2\times 10^{-4},2\times 10^{-2}]$. For person re-identification unlearning task, original model was trained over $60$ epochs with learning rate as $3\times 10^{-4}$. In the case of IMU, it is trained for $1.5\times 10^{-3}$ iterations with a learning rate of $1.5\times 10^{-3}$, with $\alpha$ set to $0.02$. For sequence modeling unlearning problem, the original model and retrained model were both trained with over $5$ epochs using softmax activation, and the learning rate was initialized at $5\times 10^{-4}$. As for SimNPO and NPO method, the value of $\beta$ was searched in interval of $[0.1,5]$ with learning rate in range $2\times 10^{-4},8\times 10^{-3}]$, while the learning rate of GA was selected from $2\times 10^{-4},10^{-2}]$. For our IMU method, learning rate was selected as $5\times 10^{-4}$ with $\alpha$ set as $0.002$. Additionally, all the unlearning methods were conducted over $50$ iterations. Lastly, in LLM unlearning scenario, $\beta$ was selected from $[1,5]$ for NPO and SimNPO, and the learning rate was searched in range $10^{-6},10^{-4}]$. Additionally, all methods in this task were executed over 10 epochs and distributed across 2 NVIDIA RTX 3090 GPUs. The official code is available at: https://github.com/goodluckisallyouneed/IMU.git.
6.2 Divergence analysis
To recap, the gradients of GA, NPO, and IMU are
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{GA}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (10)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{NPO}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[{\mathcal{W}}_{{\bm{\theta}}}({\bm{x}},{\bm{y}})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (11)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{IMU}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}},{\mathcal{D}}_{f})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right].$ (12)
@@ 6.2 Divergence analysis
Now, let us take the logistic regression as an example, where $P(y=1|x)=\sigma({\bm{z}}),P(y=0|x)=1-\sigma({\bm{z}})$ with ${\bm{z}}={\bm{\theta}}^{\top}{\bm{x}}+b,\sigma({\bm{z}})=\frac{1}{1+\exp^{-z}}$ and $b$ is a constant. The logistic model then can be written as $f_{{\bm{\theta}}}({\bm{x}})=\sigma\left(\left(2{\bm{y}}-1\right)\cdot{\bm{z}}\right)$ and the loss function would be $\ell({\bm{x}},{\bm{y}};{\bm{\theta}})=-\log f_{{\bm{\theta}}}({\bm{x}})=-\log\sigma\left(\left(2{\bm{y}}-1\right)\cdot{\bm{z}}\right)$. Hence, we have
$\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})={\bm{x}}\left(2{\bm{y}}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}\right)\right),$ (13)
To perform the gradient update for one step, we can obtain
${\bm{\theta}}^{t+1}={\bm{\theta}}^{t}+\eta\cdot{\bm{x}}\left(2{\bm{y}}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}\right)\right).$ (14)
Denote $n_{f}$ as the number of training data points ${\bm{X}}=\begin{bmatrix}{\bm{x}}_{1},{\bm{x}}_{2},\cdots,{\bm{x}}_{n_{f}}\end{bmatrix}^{\top}$. Assuming every step stochastically selects samples to update and update with stepsize $\eta$ for $t$ iterations.
For GA, we have
${\bm{\theta}}^{t}={\bm{\theta}}^{0}+\eta\sum_{i=1}^{n_{f}}\gamma_{i}^{t}\cdot{\bm{x}}_{i}\left(2{\bm{y}}_{i}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right),$ (15)
where $\gamma_{i}^{t}$ denotes the frequency of the sample ${\bm{x}}_{i}$ in the previous $t$ steps. Then, the weighted norm would be
$\|{\bm{\theta}}^{t}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}$ $=\left({\bm{\theta}}^{t}-{\bm{\theta}}^{0}\right)^{\top}{\bm{X}}^{\top}{\bm{X}}\left({\bm{\theta}}^{t}-{\bm{\theta}}^{0}\right)$
$=\eta^{2}\sum_{i=1}^{n_{f}}\sum_{j=1}^{n_{f}}\gamma_{i}^{t}\gamma_{j}^{t}\cdot\left(2{\bm{y}}_{i}-1\right)\left(2{\bm{y}}_{j}-1\right)\cdot\left(1-f_{{\bm{\theta}}}({\bm{x}}_{i})\right)\left(1-f_{{\bm{\theta}}}({\bm{x}}_{j})\right)\cdot{\bm{x}}_{i}^{\top}{\bm{X}}^{\top}{\bm{X}}{\bm{x}}_{j},$
$={\bm{a}}^{\top}{\bm{G}}{\bm{a}},$ (16)
where $a_{i}=\eta\gamma_{i}^{t}\left(2{\bm{y}}_{i}-1\right)(1-f_{{\bm{\theta}}}(x))$ so ${\bm{a}}=\begin{bmatrix}a_{1},\cdots,a_{n_{f}}\end{bmatrix}^{\top}$, and ${\bm{G}}_{ij}={\bm{x}}_{i}^{\top}{\bm{X}}^{\top}{\bm{X}}{\bm{x}}_{j}$.
Namely, for GA, we have
$\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}={\bm{a}}^{\top}{\bm{G}}{\bm{a}}.$ (17)
For NPO, we have
${\bm{\theta}}^{t}_{\texttt{NPO}}={\bm{\theta}}^{0}+\eta\sum_{i=1}^{n_{f}}\gamma_{i}^{t}\cdot{\mathcal{W}}_{{\bm{\theta}}}({\bm{x}}_{i},{\bm{y}}_{i})\cdot{\bm{x}}_{i}\left(2{\bm{y}}_{i}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right),$ (18)
@@
$\|{\bm{\theta}}^{t}_{\texttt{NPO}}$- ${\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}=({\mathcal{W}}{\bm{a}})^{\top}{\bm{G}}({\mathcal{W}}{\bm{a}}),$ (19)
where ${\mathcal{W}}=\operatorname{diag}\left({\mathcal{W}}({\bm{x}}_{1},{\bm{y}}_{1}),\cdots,{\mathcal{W}}({\bm{x}}_{n_{f}},{\bm{y}}_{n_{f}})\right)$ and ${\mathcal{W}}({\bm{x}}_{i},{\bm{y}}_{i})\in[0,2]$.
Similarly, for IMU, we have
${\bm{\theta}}^{t}_{\texttt{IMU}}={\bm{\theta}}^{0}+\eta\sum_{i=1}^{n_{f}}\gamma_{i}^{t}\cdot\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\bm{X}})\cdot{\bm{x}}_{i}\left(2{\bm{y}}_{i}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right),$ (20)

### v3

@@ 5. Conclusion and Limitations
In this paper, we introduce IMU, a novel retain-data-free MU method that dynamically allocates optimization priority to each data point according to its estimated influence scores. By leveraging a computationally efficient classifier-level influence approximation, IMU bypasses the prohibitive overhead of full-model Hessian inversion while maintaining high fidelity to the true influence signals. Extensive experiments across diverse unlearning scenarios in both vision and language tasks demonstrate that IMU effectively strikes the balance between forget quality and model utility, consistently outperforming existing baselines.
Limitation. While IMU requires computing influence scores for each forgetting data point, the cost scales with the size of ${\mathcal{D}}_{f}$. A natural mitigation is to estimate influence at the mini-batch level; however, this may introduce additional approximation error and is left for future investigation.
While IMU’s default head-only update prioritizes efficiency, it may leave frozen backbone representations partially vulnerable to sophisticated adversarial probing. However, we emphasize that the influence-guided reweighting paradigm at the core of IMU is inherently strategy-agnostic. It can be seamlessly integrated with more intensive update schemes, such as full-parameter fine-tuning or adaptive layer unfreezing (see § B.2 in the appendix), depending on the specific requirements for unlearning depth and defensive robustness. Exploring these hybrid extensions under adversarial settings constitutes a promising direction for future research.
@@
Appendix A Proofs
A.1. Classifier-Level Influence Approximation
Theorem 3.4(Classifier-Level Loss-Influence Approximation). Let ${\mathcal{L}}_{f}({\bm{\theta}})\triangleq\ell(f_{\bm{\theta}}({\mathcal{D}}_{f}))$, and let the Hessian of ${\mathcal{L}}_{f}$ at $\bm{\theta}$$^{*}$ be
${\bm{H}}^{f}_{{\bm{\theta}}^{*}}=\begin{bmatrix}{\bm{H}}_{ee}&{\bm{H}}_{ec}\\ {\bm{H}}_{ce}&{\bm{H}}_{cc}\end{bmatrix}$.
Define
(17) $\Delta^{\mathrm{full}}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})$ $\triangleq\frac{1}{N}\nabla_{{\bm{\theta}}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})^{\top}({\bm{H}}^{f}_{{\bm{\theta}}^{*}})^{-1}\nabla_{{\bm{\theta}}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}})),$
(18) $\Delta^{c}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})$ $\triangleq\frac{1}{N}\nabla_{{\bm{\theta}}_{c}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})^{\top}{\bm{H}}_{cc}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}})).$
Under 1, we have
$\Big|\Delta^{\mathrm{full}}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})-\Delta^{c}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})\Big|=\mathcal{O}\!\Bigg(\frac{1}{N}\Big($ $\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|+\|\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})\|\,\|{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|$
(19) $+\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\|\,\|\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))\|+\|\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})\|\,\|\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))\|\Big)\Bigg).$
Proof.
Let
$\mathbf{u}=\begin{bmatrix}\mathbf{u}_{e}\\ \mathbf{u}_{c}\end{bmatrix}=\nabla_{{\bm{\theta}}}{\mathcal{L}}_{f}({\bm{\theta}}^{*}),\qquad\mathbf{v}=\begin{bmatrix}\mathbf{v}_{e}\\ \mathbf{v}_{c}\end{bmatrix}=\nabla_{{\bm{\theta}}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}})),$
where $\mathbf{u}_{e}=\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})$, $\mathbf{u}_{c}=\nabla_{{\bm{\theta}}_{c}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})$, $\mathbf{v}_{e}=\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))$, and $\mathbf{v}_{c}=\nabla_{{\bm{\theta}}_{c}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))$. Define
${\bm{E}}\triangleq{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec},\qquad{\bm{S}}\triangleq{\bm{H}}_{cc}-{\bm{E}}.$
By the block inverse formula,
$({\bm{H}}^{f}_{{\bm{\theta}}^{*}})^{-1}=\begin{bmatrix}{\bm{H}}_{ee}^{-1}+{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}&-{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}\\[3.0pt] -{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}&{\bm{S}}^{-1}\end{bmatrix}$.
Under 1, $\|{\bm{H}}_{cc}^{-1}{\bm{E}}\|\ll 1$, and thus the Neumann expansion gives
${\bm{S}}^{-1}=({\bm{H}}_{cc}-{\bm{E}})^{-1}={\bm{H}}_{cc}^{-1}+{\bm{H}}_{cc}^{-1}{\bm{E}}{\bm{H}}_{cc}^{-1}+\mathcal{O}\!\left(\|{\bm{H}}_{cc}^{-1}{\bm{E}}\|^{2}\right),$
which implies
${\bm{S}}^{-1}={\bm{H}}_{cc}^{-1}+\mathcal{O}(\|{\bm{E}}\|).$
Now expand the bilinear form:
$\mathbf{u}^{\top}({\bm{H}}^{f}_{{\bm{\theta}}^{*}})^{-1}\mathbf{v}=$ $\mathbf{u}_{c}^{\top}{\bm{S}}^{-1}\mathbf{v}_{c}-\mathbf{u}_{e}^{\top}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}\mathbf{v}_{c}$
(20) $-\mathbf{u}_{c}^{\top}{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\mathbf{v}_{e}+\mathbf{u}_{e}^{\top}\left({\bm{H}}_{ee}^{-1}+{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\right)\mathbf{v}_{e}.$
Therefore,
$\mathbf{u}^{\top}({\bm{H}}^{f}_{{\bm{\theta}}^{*}})^{-1}\mathbf{v}-\mathbf{u}_{c}^{\top}{\bm{H}}_{cc}^{-1}\mathbf{v}_{c}=$ $\mathbf{u}_{c}^{\top}({\bm{S}}^{-1}-{\bm{H}}_{cc}^{-1})\mathbf{v}_{c}$
$-\mathbf{u}_{e}^{\top}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}\mathbf{v}_{c}$
$-\mathbf{u}_{c}^{\top}{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\mathbf{v}_{e}$
(21) $+\mathbf{u}_{e}^{\top}\left({\bm{H}}_{ee}^{-1}+{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\right)\mathbf{v}_{e}.$
For the first term, using ${\bm{S}}^{-1}={\bm{H}}_{cc}^{-1}+\mathcal{O}(\|{\bm{E}}\|)$, we have
$\mathbf{u}_{c}^{\top}({\bm{S}}^{-1}-{\bm{H}}_{cc}^{-1})\mathbf{v}_{c}=\mathcal{O}(\|{\bm{E}}\|)=\mathcal{O}\!\left(\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|\right).$
For the second term,
$\left|\mathbf{u}_{e}^{\top}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}\mathbf{v}_{c}\right|\leq\|\mathbf{u}_{e}\|\,\|{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|\,\|{\bm{S}}^{-1}\|\,\|\mathbf{v}_{c}\|=\mathcal{O}\!\left(\|\mathbf{u}_{e}\|\,\|{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|\right).$
For the third term,
$\left|\mathbf{u}_{c}^{\top}{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\mathbf{v}_{e}\right|\leq\|\mathbf{u}_{c}\|\,\|{\bm{S}}^{-1}\|\,\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\|\,\|\mathbf{v}_{e}\|=\mathcal{O}\!\left(\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\|\,\|\mathbf{v}_{e}\|\right).$
For the feature-interaction term,
$\mathbf{u}_{e}^{\top}\left({\bm{H}}_{ee}^{-1}+{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}{\bm{S}}^{-1}{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\right)\mathbf{v}_{e}=\mathcal{O}\!\left(\|\mathbf{u}_{e}\|\,\|\mathbf{v}_{e}\|\right).$
Substituting back $\mathbf{u}=\nabla_{{\bm{\theta}}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})$ and $\mathbf{v}=\nabla_{{\bm{\theta}}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))$ we obtain
$\Big|\Delta^{\mathrm{full}}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})-\Delta^{c}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})\Big|=\mathcal{O}\!\Bigg(\frac{1}{N}\Big($ $\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|+\|\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})\|\,\|{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|$
(22) $+\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\|\,\|\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))\|+\|\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})\|\,\|\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))\|\Big)\Bigg).$
A.2. Divergence analysis
Theorem 3.5 (Divergence Analysis). Consider logistic regression with labels $y\in\{0,1\}$ and the observed-label likelihood $\pi_{{\bm{\theta}}}(y\mid{\bm{x}})\triangleq\sigma\big((2y-1)z\big)$ where $z={\bm{\theta}}^{\top}{\bm{x}}+b$ and $\sigma(u)=\frac{1}{1+\exp(-u)}$. Assume the model is trained on $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ and denote ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, ${\bm{K}}={\bm{X}}{\bm{X}}^{\top}\in\mathbb{R}^{n\times n}$ be the Gram matrix. Then, after $t$ SGD updates with step size $\eta$, the following divergence bounds hold:
$\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2},$
$\|{\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2}\|{\bm{W}}_{\texttt{NPO}}\|^{2},$
(23) $\|{\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2}\|{\bm{W}}_{\texttt{IMU}}\|^{2}.$
where $\lambda_{\max}({\bm{K}})$ denote the largest eigenvalue of $\bm{K}$}, and ${\bm{W}}_{\texttt{NPO}}\triangleq\mathrm{diag}(w_{\texttt{NPO},1},\ldots,w_{\texttt{NPO},n}),{\bm{W}}_{\texttt{IMU}}\triangleq\mathrm{diag}(\hat{w}_{1},\ldots,\hat{w}_{n})$.
Proof.
To recap, the per-sample update directions of GA, NPO, and IMU are
(24) ${\bm{g}}_{\texttt{GA}}=-\nabla_{{\bm{\theta}}}\ell({\bm{x}}_{i},y_{i};{\bm{\theta}}),$
(25) ${\bm{g}}_{\texttt{NPO}}=-w_{\texttt{NPO},i}\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}}_{i},y_{i};{\bm{\theta}}),$
(26) ${\bm{g}}_{\texttt{IMU}}=-\hat{w}_{i}\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}}_{i},y_{i};{\bm{\theta}}).$
@@ A.2. Divergence analysis
Now, let us take the logistic regression as an example, the loss function would be $\ell({\bm{x}},y;{\bm{\theta}})=-\log\pi_{{\bm{\theta}}}(y\mid{\bm{x}})=-\log\sigma\big((2y-1)z\big)$. Hence, we have
(27) $\nabla_{{\bm{\theta}}}\ell({\bm{x}},y;{\bm{\theta}})=-\left(2y-1\right)\left(1-\pi_{{\bm{\theta}}}(y\mid{\bm{x}})\right){\bm{x}},$
and the gradient update for one step would be
(28) ${\bm{\theta}}^{t+1}={\bm{\theta}}^{t}-\eta{\bm{g}}={\bm{\theta}}^{t}-\eta\cdot\left(2y-1\right)\left(1-\pi_{{\bm{\theta}}^{t}}(y\mid{\bm{x}})\right){\bm{x}}.$
For GA, aggregating the first $t$ iterations yields
(29) ${\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}$ $=-\eta\sum_{k=0}^{t-1}(2y_{i_{k}}-1)\Big(1-\pi_{{\bm{\theta}}_{\texttt{GA}}^{k}}(y_{i_{k}}\mid{\bm{x}}_{i_{k}})\Big){\bm{x}}_{i_{k}}=-\eta{\bm{X}}^{\top}{\bm{s}}^{(t)}_{\texttt{GA}}$,
where $i_{k}$ denotes the index of the sample selected at iteration $k$ and ${\bm{s}}_{\texttt{GA}}^{(t)}\in\mathbb{R}^{n}$ is defined as
(30) $s_{\texttt{GA},i}^{(t)}\triangleq\sum_{k=0}^{t-1}{\bm{1}}[i_{k}=i]\,(2y_{i}-1)\Big(1-\pi_{{\bm{\theta}}_{\texttt{GA}}^{k}}(y_{i}\mid{\bm{x}}_{i})\Big).$
Let ${\bm{K}}\triangleq{\bm{X}}{\bm{X}}^{\top}\in\mathbb{R}^{n\times n}$ be the Gram matrix. Then, the weighted divergence satisfies
$\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}$ $=({\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0})^{\top}{\bm{X}}^{\top}{\bm{X}}({\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0})$
$=\eta^{2}({\bm{s}}_{\texttt{GA}}^{(t)})^{\top}{\bm{X}}{\bm{X}}^{\top}{\bm{X}}{\bm{X}}^{\top}{\bm{s}}_{\texttt{GA}}^{(t)}$
(31) $=\eta^{2}({\bm{s}}_{\texttt{GA}}^{(t)})^{\top}{\bm{K}}^{2}{\bm{s}}_{\texttt{GA}}^{(t)}.$
Since $\bm{K}$} is symmetric positive semidefinite, it follows that
(32) $\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}\,\|{\bm{s}}_{\texttt{GA}}^{(t)}\|^{2},$
where $\lambda_{\max}({\bm{K}})$ denote the largest eigenvalue of $\bm{K}$}.
Similarly, for NPO, aggregating the first $t$ iterations yields
(33) ${\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}$ $=-\eta\sum_{k=0}^{t-1}w_{\texttt{NPO},i_{k}}(2y_{i_{k}}-1)\Big(1-\pi_{{\bm{\theta}}_{\texttt{NPO}}^{k}}(y_{i_{k}}\mid{\bm{x}}_{i_{k}})\Big){\bm{x}}_{i_{k}}=-\eta{\bm{X}}^{\top}{\bm{W}}_{\texttt{NPO}}{\bm{s}}_{\texttt{NPO}}^{(t)},$
where ${\bm{W}}_{\texttt{NPO}}\triangleq\mathrm{diag}(w_{\texttt{NPO},1},\ldots,w_{\texttt{NPO},n})$. For analytical tractability, we treat $w_{\texttt{NPO},i}$ as a fixed per-sample weight and omit its dependence on ${\bm{\theta}}^{k}$. Then, the weighted divergence for NPO satisfies
$\|{\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}$ $=({\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0})^{\top}{\bm{X}}^{\top}{\bm{X}}({\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0})$
$=\eta^{2}({\bm{s}}_{\texttt{NPO}}^{(t)})^{\top}{\bm{W}}_{\texttt{NPO}}^{\top}{\bm{X}}{\bm{X}}^{\top}{\bm{X}}{\bm{X}}^{\top}{\bm{W}}_{\texttt{NPO}}{\bm{s}}_{\texttt{NPO}}^{(t)}$
(34) $=\eta^{2}({\bm{s}}_{\texttt{NPO}}^{(t)})^{\top}{\bm{W}}_{\texttt{NPO}}^{\top}{\bm{K}}^{2}{\bm{W}}_{\texttt{NPO}}{\bm{s}}_{\texttt{NPO}}^{(t)},$
@@
(35) $\|{\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}\,\|{\bm{W}}_{\texttt{NPO}}{\bm{s}}_{\texttt{NPO}}^{(t)}\|^{2}.$
Similarly, for IMU, aggregating the first $t$ iterations yields
(36) ${\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}$ $=-\eta\sum_{k=0}^{t-1}\hat{w}_{i_{k}}\,(2y_{i_{k}}-1)\Big(1-\pi_{{\bm{\theta}}^{k}_{\texttt{IMU}}}(y_{i_{k}}\mid{\bm{x}}_{i_{k}})\Big){\bm{x}}_{i_{k}}=-\eta{\bm{X}}^{\top}{\bm{W}}_{\texttt{IMU}}{\bm{s}}_{\texttt{IMU}}^{(t)},$
where ${\bm{W}}_{\texttt{IMU}}\triangleq\mathrm{diag}(\hat{w}_{1},\ldots,\hat{w}_{n})$. For analytical tractability, we treat $\hat{w}_{i}$ as a fixed per-sample weight and omit its dependence on ${\bm{\theta}}^{k}$, and in our implementation, $\hat{w}_{i}$ is computed from the initial model and kept fixed during the subsequent unlearning updates by default. Then, the weighted divergence for IMU satisfies
(37) $\|{\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}$ $=\eta^{2}({\bm{s}}_{\texttt{IMU}}^{(t)})^{\top}{\bm{W}}_{\texttt{IMU}}^{\top}{\bm{K}}^{2}{\bm{W}}_{\texttt{IMU}}{\bm{s}}_{\texttt{IMU}}^{(t)},$